Can someone help me on this problem?
The sum of two numbers is 33. Their difference is 11. What are the two numbers?
Thanks,
Ashley

Answer by checkley71(8403)

X+Y=33
X-Y=11
X=11+Y
(11+Y)+Y=33
11+Y+Y=33
2Y=22-11
X=11 SOLUTION
THUS
11+Y=33
Y=33-11
Y=22 SOLUTION